OpenStudy (drayder97):

a sample of air occupies 2.50 L at a temperature of 22.0 Celsius. What volume will this sample occupy inside a hot air balloon at a temperature of 43.0 Celsius. Assume that the pressure inside the balloon remains constant.

thomaster (thomaster):

Use Charle's gas law \(\LARGE \dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}\) T= temperature in Kelvin V= volume If you would put your date in the equation, you'd get this: \(\LARGE \dfrac{2.5}{295.15}=\dfrac{V_2}{316.15}\) \(\LARGE 8.47*10^{-3}=\dfrac{V_2}{316.15}\) \(\LARGE V_2=8.47*10^{-3}*316.15\)